OpenStudy (anonymous):

Anna asks her broker to divide her investment of $2000 among three funds, X, Y, and Z. She decides that her investment in X should be twice her investment in Z. During the first quarter, X earned 4.5%, Y earned 2.6%, and Z lost 0.2%. Anna earned $58 on her investment during that period. How many dollars did the broker put into each fund?

OpenStudy (anonymous):

\[ x+y+z = 2000 \]\[ 0.045x+0.026y-0.002z = 58 \]

OpenStudy (anonymous):

\[ x = 2z \]

OpenStudy (anonymous):

This is a matrix: \[ \begin{bmatrix} 1&1&1\\ 0.045&0.026&0.002 \\ 1&0&-2 \end{bmatrix} \begin{bmatrix} x\\y\\z \end{bmatrix} = \begin{bmatrix} 2000\\58\\0 \end{bmatrix} \]